# Quantum Entanglement of Anyonic Charges and Emergent Spacetime Geometry

### v1

@@ Appendix A Gaussian Curvature of the Induced Metric
which explicitly encodes how distances along the $x$-direction vary with position in $y$. has Gaussian curvature given by
@@
In our case,
$f(y)=\frac{W}{y_{0}\cosh^{2}(y/y_{0})}.$ (17)
@@ Appendix B Einstein Field Equations
We first find the Riemann curvature tensor from the metric [Susskind2023]. Consider a metric of the following form:
@@
The affine connection [zee2013einstein] is given by
@@
Lastly, if $\lambda=y$, $\sigma=\lambda=y$. The last term of Eq. (19) gives $f^{\prime}(y)f(y)$ for $\mu=\nu=x$, while the first two terms give vanishing result since all metric components involving $y$ index are constants.
@@
$R^{\lambda}{}_{\mu\nu\sigma}=\partial_{\nu}\Gamma^{\lambda}_{\mu\sigma}-\partial_{\sigma}\Gamma^{\lambda}_{\mu\nu}+\Gamma^{\eta}_{\mu\sigma}\Gamma^{\lambda}_{\eta\nu}-\Gamma^{\eta}_{\mu\nu}\Gamma^{\lambda}_{\eta\sigma}$ (21)
The anti-symmetry under $\nu\longleftrightarrow\sigma$ is manifest.
@@ Appendix C Correlations in the Hyperbolic Space
A key feature of conformal field theories (CFTs) in flat (Euclidean) space is the algebraic decay of two-point correlators [Cardybook]:
$\langle\mathcal{O}(x),\mathcal{O}(0)\rangle\sim\frac{1}{|x|^{2\alpha}}.$ (28)
In the AdS/CFT correspondence, the boundary theory is conformally invariant, so this same functional form arises [Witten1998, Hartnoll2009]. In the AdS bulk, by contrast, the geodesic distance between two boundary points grows only logarithmically with their coordinate separation at large distances:
$d_{\mathrm{hyp}}(x_{1},x_{2})\sim\log\left(\frac{|x_{1}-x_{2}|}{\epsilon}\right),$ (29)
@@
$\langle\mathcal{O}(x_{1}),\mathcal{O}(x_{2})\rangle\sim e^{-\alpha d_{\mathrm{hyp}}(x_{1},x_{2})}.$ (30)

### v2

@@ Appendix A Gaussian Curvature of the Induced Metric
which explicitly encodes how distances along the $x$-direction vary with position in $y$. The metric has Gaussian curvature given by
@@
In our case, $f(y)=W/[y_{0}\cosh^{2}(y/y_{0})]$ so the curvature takes the following form
$K=\frac{2}{y_{0}^{2}}\left(\frac{3}{\cosh^{2}{(y/y_{0})}}-2\right).$ (19)
@@ Appendix B Einstein Field Equations
We first find the Riemann curvature tensor from the metric [50]. Consider a metric of the following form:
@@
The affine connection [61] is given by
@@
Lastly, if $\lambda=y$, $\sigma=\lambda=y$. The last term of Eq. (21) gives $f^{\prime}(y)f(y)$ for $\mu=\nu=x$, while the first two terms give vanishing result since all metric components involving $y$ index are constants. We have
@@
$R^{\lambda}{}_{\mu\nu\sigma}=\partial_{\nu}\Gamma^{\lambda}_{\mu\sigma}-\partial_{\sigma}\Gamma^{\lambda}_{\mu\nu}+\Gamma^{\eta}_{\mu\sigma}\Gamma^{\lambda}_{\eta\nu}-\Gamma^{\eta}_{\mu\nu}\Gamma^{\lambda}_{\eta\sigma},$ (23)
where the anti-symmetry under $\nu\longleftrightarrow\sigma$ is manifest.
@@ Appendix C Correlations in the Hyperbolic Space
A key feature of conformal field theories (CFTs) describing critical systems in flat (Euclidean) space is the algebraic decay of two-point correlators [4]:
$\langle\mathcal{O}(x)\mathcal{O}(0)\rangle\sim\frac{1}{|x|^{2\alpha}}.$ (30)
In the AdS/CFT correspondence, the boundary theory is conformally invariant, so this same functional form arises [59, 18]. In the AdS bulk, by contrast, the geodesic distance between two boundary points grows only logarithmically with their coordinate separation at large distances:
$d_{\mathrm{geo}}(x_{1},x_{2})\sim\log\left(\frac{|x_{1}-x_{2}|}{\epsilon}\right),$ (31)
@@
$\langle\mathcal{O}(x_{1}),\mathcal{O}(x_{2})\rangle\sim e^{-\alpha d_{\mathrm{geo}}(x_{1},x_{2})}.$ (32)